OpenStudy (lgbasallote):

what are the values of R again? 0.0821 is one. what is the other one?

mathslover (mathslover):

8.3144621 J/K per mole

mathslover (mathslover):

mentioned by you : R =0.08205746 L atm K−1 mol−1

mathslover (mathslover):

@lgbasallote got it?

mathslover (mathslover):

PV = nRT R = universal gas constant

OpenStudy (lgbasallote):

wait what is the unit of that 8.31 thingy?

mathslover (mathslover):

J/ K per mole : \(\large{J K^{-1} mol^{-1}}\)

OpenStudy (lgbasallote):

\[\huge \frac{J}{K \cdot \text{mol}}\] that?

mathslover (mathslover):

Yes . \[\large{\frac{J}{mol K}}\]

OpenStudy (lgbasallote):

when is it used?

mathslover (mathslover):

PV = nRT

mathslover (mathslover):

Using the value of R we can calculate the rest variables..

OpenStudy (lgbasallote):

i meant when is 0.0821 used and when is 8.31 used?

OpenStudy (lgbasallote):

you there @mathslover ?

mathslover (mathslover):

Oh sorry i was there away for some time .. actually 0.0821 and 8.31 ... are both R but have diff. units They are both used in PV=nRT

mathslover (mathslover):

actually R has several values dependent on : units

mathslover (mathslover):

these are the values

OpenStudy (lgbasallote):

hmm i just need to know these two and when they're used

OpenStudy (lgbasallote):

do you kknow when?

mathslover (mathslover):

they are both used in PV=nRT formula dependent on the units .. for example we have units of RT something similar to 0.0821 unit .. then we use this and if the units are like J k^-1 mol^-1 then we use 8.31

OpenStudy (lgbasallote):

huh?

OpenStudy (lgbasallote):

isnt there a specific sceneraio?

OpenStudy (lgbasallote):

scenario*

mathslover (mathslover):

yes there is : but available on wikipedia ..

mathslover (mathslover):

Well i can give you summary : 1) R has several values .. for example 1 m = 100 cm = 1000 mm and so o n similarly R has several values dependent on units 2) Use of 0.081 and 8.31 --> Both can be used in the formula : PV=nRT it depends on units given for example : we generally use 2 second to calculate speed if distance =4 m . similarly here .. we will use that value which satisfies result

mathslover (mathslover):

I hope that it helps you

OpenStudy (lgbasallote):

hmm okay...i guess i'll just fully understand when i see it in action

mathslover (mathslover):

.. hmn yes .. why not do some practice problems on that?

mathslover (mathslover):

best of luck